OpenStudy (anonymous):

Suppose that r varies directly with s and inversely with t, and r = 2 when s = 3 and t = 12. What is the value of r when s = 5 and t = 4? 2/5 32/5 10 20

OpenStudy (anonymous):

2 * (3/2) = 3 s 2 * 6 = 12 t 3 (5/3) = 5 so 2(5/3) = 10/3 [s] <--- Directly 12 (1/3) = 4 so (10/3)(3/1) = 10[s] <--- Inverse Answer is 10.